OpenStudy (anonymous):

simplify the trig expression....csc^3x-csc^2x-cscx+1

OpenStudy (anonymous):

i dont know how to factor it

jimthompson5910 (jim_thompson5910):

are you able to factor something like z^3 - z^2 - z + 1 ?

OpenStudy (lochana):

plug cosx = 1

OpenStudy (anonymous):

i think i can factor it.

OpenStudy (lochana):

of course now you can. we found that (cosx-1) is a factor

OpenStudy (anonymous):

where is cosx-1 coming from

OpenStudy (lochana):

find A,B and C. then you can refactor it

jimthompson5910 (jim_thompson5910):

@lochana those are cosecants, not cosines

OpenStudy (lochana):

yes. that's right. I didn't see it clearly. please change cosx into cscx. sorry:(

OpenStudy (lochana):

so plug cscx = 1 now you can say \( csc^3x - csc^2x - cscx + 1 = (cscx -1)(Acsc^2x + Bcscx + C )\)

OpenStudy (anonymous):

why am i plugging cscx=1??

OpenStudy (lochana):

because your equation becomes 0 when cscx = 1.

OpenStudy (lochana):

0 means it's a factor. only factors can do that.

OpenStudy (anonymous):

im confused

OpenStudy (lochana):

familar with \(x^2 + 4x + 4\)?

OpenStudy (anonymous):

no

OpenStudy (lochana):

\(x^2+4x+4 = (x+2)^2\) right?

OpenStudy (anonymous):

yes

OpenStudy (lochana):

now plug x = -2 in \((x+2)^2\)

OpenStudy (lochana):

what happened?

OpenStudy (lochana):

it became 0. yes? therefore when you plug x = -2 in \(x^2 + 4x +4\), it becomes 0 since \((x+2)^2 = x^2+4x+4\). same way, \(csc^3x - csc^2x - cscx +1 \) becomes 0 when \(cscx = 1\) means (cscx-1) is a factor

OpenStudy (lochana):

and A = 1; B = 0; C = -1

OpenStudy (lochana):

so now it is\((cscx-1)(csc^2x - 1 )\), well now it is not that hard. answer will be \((cscx-1)(cscx+1)(cscx -1 )\)